At 10Am, adam started out at 60 kilometers per hour to catch tom , who had a three hour head start. If adam was still 70 kilometers behind at 6 pm, how fast was tom driving?
Found 2 solutions by Edwin McCravy, stanbon:
Answer by Edwin McCravy(20086)

At 10Am, adam started out at 60 kilometers per hour to catch tom , who had a
three hour head start. If adam was still 70 kilometers behind at 6 pm, how fast
was tom driving?

Make this chart
        
         DISTANCE  RATE  TIME
Adam                                          
Tom                            

Let Tom's rate be x.  Fill this in and 
also Adam's rate, which is 60.

         DISTANCE  RATE  TIME
Adam                60                        
Tom                  x      

At 6 pm, Adam had driven since 10AM, 
which is 8 hours.  Tom has a three hour 
head start, so at 6 PM Tom had driven 
for 11 hours. So fill in Adam's and 
Tom's times:

         DISTANCE  RATE  TIME
Adam                60     8                  
Tom                  x    11

Use DISTANCE = RATE × TIME to fill in 
the distances:

         DISTANCE  RATE  TIME
Adam       480      60     8                  
Tom        11x       x    11

Since Adam was 70 km behind Tom at 
6 PM, then

Adam's distance = Tom's distance minus 70

            480 = 11x - 70

Solve that and you'll get that Tom's rate 
is 50 km/h

Edwin

Answer by stanbon(75887)

At 10Am, adam started out at 60 kilometers per hour to catch tom , who had a three hour head start. If adam was still 70 kilometers behind at 6 pm, how fast was tom driving?
--------------
Adam DATA:
rate= 60 km/hr ; time = 8 hr. ; distance = tr= 480 km
---------
Tom DATA:
time=8+3 = 11 hr;
Let Tom's rate be x km/hr
In 3 hrs he goes 3x km
He then covers the 480 km that Adam covers later
He then is 70 km ahead of Adam
Tom's distance = 3x+480+70=3x+550 km
Tom's distance is also 11x because he was on the road for 11 hours
EQUATION:
3x+550=11x
8x=550
x=68.75 km/hr (This is Tom's rate)
Cheers,
Stan H.
rate=d/t=550/11=50 km/hr
-----------------
Tom's rate is 50 km/hr.